The demand for product X depends on the price of product X as well as the average household income (Y) according to the following relationship

Qdx = 750 - 25 P + 0.001Y

The supply of product X is positively related to own price of product X and negatively dependent upon W, the price of some input. This relationship is expressed as:

Qsx = 60 + 15 P - 2 W
Given that Y = 40,000 and W = 6, what is the:
1. Equilibrium price? ____________
2. Equilibrium quantity? ____________
Suppose that income increases to 50,000 and W remains constant. What is the new:
3. Equilibrium price? ____________
4. Equilibrium quantity? ____________
Assuming that income remains constant at 50,000 and W increases to 11, what is the new:

5: Equilibrium price? ____________
6. Equilibrium quantity? ____________
1
Expert's answer
2020-04-06T09:58:50-0400

1. "Y=40000; W=6" "750-25P+0.001\u00d740000=60+15P-2\u00d76;"

"750-25P+40=60+15P-12;"

"P=18.55;"

2. "Qd=750-25\u00d718.55+0.001\u00d740000; Qd=750-463.75+40;"

"Qd=326.25;"

"Qs=60+15\u00d718.55-2\u00d76; Qs=60+278.25-12;"

"Qs=326.25"

3.

"Y=50000; W=6;"

"750-25P+0.001\u00d750000=60+15P-2\u00d76;"

"750-25P+50=60+15P-12;\n800-48=25P+15P;"

"40P=752; P=18.8;"

4. "Qd=750-25\u00d718.8+0.001\u00d750000;"

"Qd=750-470+50; Qd=330;"

"Qs=60+15\u00d718.8-2\u00d76; Qs=60+282-12;"

"Qs=330;"

5.

"Y=50000; W=11;"

"750-25P+0.001\u00d750000=60+15P-2\u00d711;"

"750+50-60+22=25P+15P;"

"40P=762; P=19.05";

6. "Qd=750-25\u00d719.05+50; Qd=750-476.25+50;"

"Qd=323.75;"

"Qs=60+15\u00d719.05-2\u00d711;"

"Qs=60+285.75-22; Qs=323.75;"